Au2S3 + 3 H2 = 2 Au + 3 H2S


Find the mass of gold produced when 2.00 g of hydrogen gas is reacted with an excess of gold (III) sulfide. 


1
Expert's answer
2021-03-01T04:42:07-0500
"n(H2)=m\/M=2\/2=1 mol"

"1\/3=n(Au)\/2"

"n(Au)=0.67 mol"

"m(Au)=0.67*197=132 g"